### v1

# Quasi-Hereditary Orderings of Nakayama Algebras

Abstract.
To determine an algebra is quasi-hereditary is a difficult problem. An effective method, Green-Schroll set, is introduced in this paper to tackle this problem. It is well known that an algebra is quasi-hereditary if and only if it admits a quasi-hereditary ordering of simple modules. Let $A$ be a Nakayama algebra. We prove a necessary and sufficient criterion to determine whether an ordering of simple modules is quasi-hereditary on $A$, and $A$ is quasi-hereditary if and only if its Green-Schroll set is nonempty. This seems to be the simplest characterization currently known, since it does not involve any algebraic concepts. A general iteration formula for the number $q(A)$ of all quasi-hereditary orderings of $A$ is given via Green-Schroll set. The $q$-ordering conjecture is proved to be true for $A$.
Supported by National Natural Science Foundation of China under Grant No. 123B1026.
* Corresponding author.
E-mail addresses: zyh@sjtu.edu.cn (Yuehui Zhang), 21-zxq@sjtu.edu.cn (Xiaoqiu Zhong)
1. Introduction
The theory of quasi-hereditary algebras has been extensively studied since its introduction by E. Cline, B. Parshall, and L. Scott in their seminal paper [1] in 1988. In this theory, the properties of a quasi-hereditary algebra $A$ are heavily dependent on a specific ordering of the isomorphism classes of simple $A$-modules $\mathcal{S}$. This ordering ensures that every indecomposable projective module can be filtered by the Weyl modules, which are constructed through a special process related to the ordering. Such an ordering is termed a quasi-hereditary ordering (or $q$-ordering) on $A$. Thus, when referring to a quasi-hereditary algebra, it typically includes specification of the $q$-ordering.
This naturally leads to the question: How many $q$-orderings does an arbitrary algebra $A$ admit? Denote this number by $q(A)$, and let $|\mathcal{S}|=n$, where $n$ is the rank of the Grothendieck group of $A$. It is evident that $0\leq q(A)\leq n!$. A classical result, established by Dlab and Ringel, asserts that $q(A)=n!$ if and only if $A$ is hereditary [2].
While obtaining an explicit formula for $q(A)$ for arbitrarily given algebra $A$ remains challenging, significant progress has been made, particularly for Nakayama algebras. In 1990, M. Uematsu and K. Yamagata demonstrated that Morita equivalence preserves quasi-heredity [3], meaning that if two algebras $A$ and $B$ are Morita equivalent, then $A$ is quasi-hereditary if and only if $B$ is. They also established that a Nakayama algebra $A$ is quasi-hereditary if and only if it has a simple module of projective dimension 2. In 2000, the first author of the present paper and Y. Li proved [4] that $q(A)\leq\dfrac{2}{3}n!$ for tree-type quasi-hereditary algebras of one generator. Furthermore, they conjectured that this inequality holds for all non-hereditary algebras, namely, they conjectured that $q(A)>\dfrac{2}{3}n!$ if and only if $q(A)=n!$ if and only if $A$ is hereditary. This conjecture, termed the ”$q$-ordering conjecture,” poses an intriguing challenge in the field. In 2008, the first author, along with L. Wu and C. Gao, provided a necessary and sufficient condition for an ordering on $\mathcal{S}$ to be a $q$-ordering for $A_{n}$-type algebras with exactly two generators. They also derived an explicit formula for $q(A)$ in this context [5], and conclusively affirmed the $q$-ordering conjecture in that specific case.
In 2019, E.L. Green and S. Schroll [6] provided a necessary and sufficient condition for the existence of a $q$-ordering on $\mathcal{S}$ for all monomial algebras using quiver language.
A directed Nakayama algebra is concave if and only if its Kupisch series first increases, then decreases. In 2022, C.M. Ringel [7] gave the classification of the concave Nakayama algebras which are higher Auslander algebras.
An artin algebra $A$ is said to be eventually periodic if the $A^{op}$-module $A$ is eventually periodic. In 2022, S. Usui [8] proved that a finite-dimensional connected Nakayama algebra is eventually periodic if and only if it is of infinite global dimension. This result may be viewed as a counterpart of quasi-hereditary algebras, since quasi-hereditary algebras are of finite global dimensions.
A noteworthy recent result was established by R. Marczinzik and E. Sen [9] in 2022. They introduced the concept of ”$S$-connected” for a finite-dimensional algebra $A$, defined by the property that the projective dimensions of simple $A$-modules form an interval. Their key finding is that a Nakayama algebra $A$ is $S$-connected if and only if $A$ is quasi-hereditary.
The primary objective of this paper is as follows. Let $A$ be a Nakayama algebra. We define its Green-Schroll set as a special subset of all simple modules. We show that $A$ is quasi-hereditary if and only if its Green-Schroll set is nonempty. This seems to be the simplest characterization currently known. We characterize simple modules of projective dimensions less than 3 via Green-Schroll set. We also give a complete classification of all $q$-orderings and provide an iteration formula for $q(A)$. This formula conclusively affirms the $q$-ordering conjecture for all Nakayama algebras. Finally, we present a counterexample to demonstrate that the $q$-ordering conjecture does not hold when $A$ is not monomial.
Unless otherwise specified, all algebras discussed in this paper are assumed to be (connected) finite-dimensional over a fixed algebraically closed field $K$, with the unit element of an algebra $A$ denoted as $1_{A}$. For any set $X$, its cardinality is denoted by $|X|$.
2. Preliminaries
2.1. Quasi-hereditary Algebra
Let $A$ be an algebra, and let $\Lambda$ be a finite poset in bijective correspondence with the set of iso-classes $\mathcal{S}$ of simple $A$-modules. The poset $\Lambda$ is referred to as the weight poset of $A$, with its elements being termed weights of $A$. For each weight $\lambda\in\Lambda$, let $S_{\lambda}$ denote the corresponding simple module and $P(\lambda)$ denote its projective cover. Define $\Delta(\lambda)$ as the maximal factor module of $P(\lambda)$ with composition factors of the form $S_{\mu}$, where $\mu\preceq\lambda$. Denote by $\Delta$ the full subcategory of $A$-mod consisting of all $\Delta(\lambda)$, where $\lambda\in\Lambda$. The modules in $\Delta$ are termed Weyl modules.
Let ${\mathscr{F}}(\Delta)$ denote the class of all $A$-modules filtered by $\Delta$, meaning there exists a filtration
$0=M_{t}\subset M_{t-1}\subset\cdots\subset M_{1}\subset M_{0}=M$
such that each factor module $M_{i-1}/M_{i}$ is isomorphic to some Weyl module, for all $1\leq i\leq t$. Modules in ${\mathscr{F}}(\Delta)$ are referred to as good modules. For simplicity, ${\mathscr{F}}(\Delta)$ is denoted simply as ${\mathscr{F}}$.
Definition 2.1.
The algebra $A$ is said to be quasi-hereditary with respect to $(\Lambda,\preceq)$ if for each $\lambda\in\Lambda$ we have:
(1) $End_{A}(\Delta(\lambda))\simeq K$;
(2) $P(\lambda)\in{\mathscr{F}}$.
By $(A,\Lambda)$ we denote a quasi-hereditary algebra $A$ with weight poset $\Lambda$. We often omit the weight poset $\Lambda$ if it is obvious in the context.
A module $M$ satisfying condition (1) of 2.1 (i.e. $End_{A}M\simeq K$) is usually called to be Schurian, so an algebra $A$ is quasi-hereditary if and only if all Weyl modules are Schurian and all projective modules are good modules.
Throughout, $\mathscr{O}(A)$ denotes the set of all $q$-orderings of $A$, thus $q(A)=|\mathscr{O}(A)|$.
2.2. Nakayama Algebras
According to [3], Proposition 2.3, Morita equivalent algebras share the same quasi-heredity properties. Additionally, it is well-known that any finite-dimensional algebra is Morita equivalent to a quotient algebra of a path algebra defined by a finite quiver. Hence, we will concentrate our investigation on the quasi-heredity of Nakayama algebras by categorizing them into directed and cyclic types. Directed Nakayama algebras are those Morita equivalent to algebras defined by directed quivers, while cyclic Nakayama algebras are those Morita equivalent to algebras defined by cyclic quivers.
2.2.1. Directed Nakayama Algebras
For any $n\geq 2$, $A_{n}$ denotes the following quiver:
Figure 1. $A_{n}$
The arrow from vertex $j$ to $j+1$ is denoted by $\alpha_{j}$, where $11\leq j\leq n-$ 1.
The oriented quiver $A_{n}$ is commonly referred to as the linearly ordered Dynkin diagram with $n$ vertices.
An algebra $A$ is termed a directed Nakayama algebra of $A_{n}$ type if $A$ is a factor algebra $KA_{n}/I$ of the quiver algebra $KA_{n}$, where $I$ is an admissible and two-sided ideal of $KA_{n}$.
2.2.2. Cyclic Nakayama Algebra
For any $n\geq 2$, $\widetilde{A}_{n}$ denotes the following quiver:
Figure 2. $\widetilde{A}_{n}$
In the cyclic case, the arrows from vertex $j$ to $j+1$ and from vertex $n$ to $1$ are denoted by $\alpha_{j}$, where $11\leq j\leq n-$ 1, and $\alpha_{n}$ respectively.
The oriented quiver $\widetilde{A}_{n}$ is commonly referred to as the Euclidean diagram with $n$ vertices.
An algebra $A$ is considered a cyclic Nakayama algebra if $A$ is a factor algebra $K\widetilde{A}_{n}/I$ of the path algebra $K\widetilde{A}_{n}$ with $I$ being an admissible and two-sided ideal of $K\widetilde{A}_{n}$.
Cyclic Nakayama algebra is also known as Euclidean type Nakayama algebra.
2.3. Notations of Simple Modules, Hooks and Denouements of Paths
Let $Q$ be any finite quiver, and let $A=KQ/I$ be a finite-dimensional algebra with $I$ being an admissible ideal of $KQ$. For simplicity, we denote the simple $A$-module $S_{x}$ corresponding to the vertex $x$ (i.e., $S_{x}=A/AeA$ with $e=1_{A}-x$) simply as $x$. Thus, the symbol $x$ simultaneously represents a number, a vertex, and a simple module. Consequently, the weight poset $\Lambda$ of $A$ can be chosen as the set $\mathcal{S}$ of all simple $A$-modules. This notation system prevents confusion when we use $\preceq$ (or its reverse order $\succeq$) to denote the partial order of the weight poset $\Lambda$.
Consider $g\in I$ to be a path of length(= the number of arrows $+1$) $\ell_{g}\geq 3$ in $KQ$. The origin vertex of $g$ is referred to as the ”hook” of $g$, denoted as $h(g)$, while the end vertex of $g$ is termed the ”denouement” of $g$, denoted as $d(g)$. The ordered multiple set $\{h(g),h(g)+1,\dots,d(g)\}$ of simple $A$-modules associated with $g$ is called the ”hood” of $g$ and denoted as $Hod(g)$. The multiple set $Hod(g)\setminus\{h(g),d(g)\}$ of simple $A$-modules is denoted by $Int(g)$, where the elements of $Int(g)$ are referred to as interior simple modules of $g$. It is noteworthy that $|Hod(g)|=\ell_{g}$ and $|Int(g)|=\ell_{g}-2$.
Given any simple ordering $\preceq$ of $\mathcal{S}$. Denote by $\max Hod(g)$ the maximum of $Hod(g)$ with respect to $\preceq$.
2.4. Definition of Green-Schroll Set for Nakayama Algebras
We will use $Q_{n}$ to denote both the quiver $A_{n}$ and $\widetilde{A}_{n}$. Thus, an arbitrarily given Nakayama algebra is Morita equivalent to $KQ_{n}/I$ for some $n$ with an admissible ideal $I$.
Since Morita equivalence preserves quasi-heredity, to analyze $q$-orderings of Nakayama algebras, it suffices to focus on Nakayama algebras of the form $KQ_{n}/I$. Each generator $g=\alpha_{d(g)-1}\alpha_{d(g)-2}\cdots\alpha_{h(g)}$ of $I$ is a path from its origin vertex $h(g)$ to its end vertex $d(g)$ of length $\ell_{g}\geq 3$, where $d(g)=h(g)+\ell_{g}-1$. In the event that a number $v$ referring to a vertex or a simple module exceeds $n$, we understand it as the vertex or simple module $v\,\text{mod}(n)$.
Note that any admissible ideal $I$ of $KQ_{n}$ is generated by a minimal set $\{g_{1},\dots,g_{k}\}$ of generators, where each $g_{i}$ is a path from its origin vertex $v_{i}$ and $v_{1}<v_{2}<\dots<v_{k}$. Throughout, when we write an admissible ideal as $I=\{g_{1},\dots,g_{k}\}$, the set $\{g_{1},\dots,g_{k}\}$ is always supposed to be minimal.
Definition 2.2.
For a Nakayama algebra $A=KQ_{n}/I$ with $I=\langle g_{1},\dots,g_{k}\rangle$. The Green-Schroll set of $A$, denoted by $\mathscr{X}_{A}$ or simply $\mathscr{X}$, is the set $\mathcal{S}-\bigcup_{i=1}^{k}\text{Int}(g_{i})$. Let $\mathscr{X}^{0}$, $\mathscr{X}^{1}$, $\mathscr{X}^{2}$, respectively, be the subset of $\mathscr{X}$ whose elements do not belong to any hood, are either hooks or denouements of some hoods but not both, and are both hooks and denouements of some hoods, respectively.
Note that $\mathscr{X}^{0}$, $\mathscr{X}^{1}$, $\mathscr{X}^{2}$ form a partition of $\mathscr{X}$, namely, they are disjoint subsets of $\mathscr{X}$ and $\mathscr{X}=\mathscr{X}^{0}\cup\mathscr{X}^{1}\cup\mathscr{X}^{2}$.
As we will see in Section 3 (Theorem 3.4), the quasi-heredity of a Nakayama algebra is completely determined by its Green-Schroll set.
2.5. Notations for indecomposable modules over Nakayama Algebras
Since Nakayama algebras are serial, each indecomposable module possesses precisely one composition series (from its top to its socle), completely determined by its top and dimension (length). To denote the unique indecomposable projective module with top $t$ and dimension $l$, we employ the symbol $(t;l)$. Hence, the indecomposable projective module $P(j)$, where $top(P(j))=j$, is exactly the module $(j;l)$, where $l=\dim_{K}P(j)$. Furthermore, from this representation, it is evident that $radP(j)=(j+1;l-1)$.
To deal with the $q$-orderings, note that the condition (1) in the Definition 2.1 is automatically satisfied for all directed algebras; for a cyclic Nakayama algebra $KQ_{n}/I$, to assure the condition (1), the indecomposable projective module covering the largest simple module (according to the corresponding $q$-ordering) must be multiplicity free, thus its length has to be at most $n$, forcing the shortest length of all generators of $I$ is at most $n$.
Another easy fact of a directed Nakayama algebra is that it has a unique simple projective module $P(n)$ and obviously $\Delta(n)=P(n)$ under all simple orderings. So, $P(n)$ always satisfies the condition (2) in the Definition 2.1.
Example 2.3.
Let $A=KA_{3}/<\alpha_{2}\alpha_{1}>$ with
Figure 3. $A_{3}$
Choose a simple ordering on the weight poset $\Lambda=\mathcal{S}$ as $1\succeq 2\succeq 3$, then $P(k)=\Delta(k)$, for $k=1,2,3$, so $1\succeq 2\succeq 3$ is a $q$-ordering. Similarly, one checks easily that $1\succeq 3\succeq 2$ is also a $q$-ordering. However, if one chooses the weight poset $\Lambda$ with ordering the simple modules as $2\succeq 1\succeq 3$, then $P(k)=\Delta(k)$, for $k=2,3$, but $\Delta(k)=1$ and $P(1)\not\in\mathscr{F}$, so $2\succeq 1\succeq 3$ is not a $q$-orderings. In fact, check other 3 orderings, $3\succeq 1\succeq 2$ and $3\succeq 2\succeq 1$ are $q$-ordering, $2\succeq 3\succeq 1$ is not. So $q(A)=4$.
Example 2.4.
Let $A=K\tilde{A}_{3}/<\alpha_{2}\alpha_{1}>$ with
Figure 4. $\tilde{A}_{3}$
Similar to Example 2.3, one can easily get $q(A)=4$. So q-conjecture holds in this case.
3. Quasi-hereditary Nakayama Algebras
In this section, we use hoods to characterize $q$-orderings over Nakayama algebras, then prove a very simple criterion for quasi-heredity of Nakayama algebras via the Green-Schroll set.
3.1. Quasi-hereditary orderings of Nakayama Algebras
Let $g$ be a path of $KQ$. If the cardinality $|Hod(g)|$ of the hood $Hod(g)$ is smaller enough, then the indecomposable projective module $P(h(g))$ is Schurian, forcing the Weyl module $\Delta(h(g))$ is also Schurian, as stated in the following lemma.
Lemma 3.1.
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},\dots,g_{k}>$. If $|Hod(g_{i})|\leq n+1$, $i=1,\dots,k$, then every Weyl module is Schurian under any simple ordering.
Next, we show the following result, called Ladder lemma.
Lemma 3.2.
Let $A=KQ_{n}/I$ be a Nakayama algebra, $1\leq k\leq n$. Given any simple ordering $\preceq$ with good module category $\mathscr{F}$.
(1) If $radP(k)\in\mathscr{F}$, then $P(k)\in\mathscr{F}$.
(2) If k is not a hook and $P(k+1)\in\mathscr{F}$, then $P(k)\in\mathscr{F}$.
(3) Let g be a generator of $I$ with hook $h(g)=k$ and denouement $d(g)$. If $d(g)=\max Hod(g)$ and $P(k)$ is Schurian, then $P(k)\in\mathscr{F}$.
Proof.
(1) If $P(k)=(k;\ell)$, then $radP(k)=(k+1;\ell-1)$. Suppose $\Delta(k)=(k;\ell^{\prime})$, $1\leq\ell^{\prime}\leq\ell$. There are 3 cases.
(i) $\ell^{\prime}=\ell$.
Then $\Delta(k)=(k;\ell)=P(k)$, thus $P(k)\in\mathscr{F}$.
(ii) $\ell^{\prime}=1$.
Then $\Delta(k)=k$. It is obvious that $P(k)\in\mathscr{F}$ since $P(K)$ is filtered by good modules $radP(k)$ and $\Delta(k)$.
(iii) $1<\ell^{\prime}<\ell$.
Thus $k=\max[k,k+\ell^{\prime}-1]$. So, $k+\ell^{\prime}\not\in\Delta(i),\forall\enspace k\leq i\leq k+\ell^{\prime}-1$ Since $radP(k+1)\in\mathscr{F}$, there exists a module $M=(k+\ell^{\prime};\ell-\ell^{\prime})\in\mathscr{F}$. Thus $P(k)\in\mathscr{F}$ as $P(k)$ is filtered by good modules $M$ and $\Delta(k)$.
(2) If $k$ is a hook, then $radP(k)=P(k+1)$. So $radP(k)\in\mathscr{F}$. By (1), $P(k)\in\mathscr{F}$.
(3) Let $d(g)=\max Hod(g)$ and $P(k)$ Schurian. Then $P(k)=(k;\ell_{g}-1)$. Since $d(g)=\max Hod(g)$, each Weyl module $\Delta(i)$ with $k\leq i\leq\ell_{g}-1$ is completely determined by the restriction of $\preceq$ on $Hod(g)$ and irrelevant to any simple modules outside $Hod(g)\setminus\{d(g)\}$. Now the full subcategory $\mathcal{C}$ of $modA$ generated by the simple modules of $Hod(g)\setminus\{d(g)\}$ is Morita equivalent to (indeed isomorphic to) the module category of the hereditary algebra $KA_{\ell_{g}-1}$, thus $\mathcal{C}$ is also a hereditary abelian category. Therefore every simple ordering on $Hod(g)\setminus\{d(g)\}$ is a $q$-ordering, forcing $P(k)\in\mathscr{F}$. ∎
Ladder lemma produces the following easy criterion for $q$-orderings.
Theorem 3.3.
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},\dots,g_{k}>$. Let $\preceq$ be a simple ordering of $A$. Then $\preceq$ is quasi-hereditary if and only if $\max Hod(g_{i})\notin Int(g_{i})$, $\forall i=1,\dots,k$.
Proof.
Sufficiency. Assume there exists a $Hod(g)$ whose maximum $o\in Int(g)$ under $\preceq$. There are following two cases.
(1) If $o=h(g)$ or $o=d(g)$. Then the multiplicity of the simple module $o$ in $\Delta(o)$ is at least 2, so $dim_{K}End_{A}(\Delta(o))\geq 2$, so the condition (1) in Definition 2.1 of quasi-hereditary algebra fails, contradicting that $\preceq$ is a $q$-ordering.
(2) If $o\neq h(g),d(g)$. $\forall i\in Hod(g)$, $o\notin\Delta(i)$ except $\Delta(o)$. By direct computation, there exists $j\geq 0$, $soc\Delta(o)=d(g)+j,socP(h(g))=d(g)-1$. Note that $o\succeq h(g),o\in P(h(g))$, so $P(h(g))$ does not admit a $\Delta$-filtration, thus $P(h(g))\notin\mathscr{F}$, so the condition (2) in Definition 2.1 fails, which contradicts that $\preceq$ is a $q$-ordering.
Necessity. Assume $\preceq$ satisfies $\max Hod(g_{i})\notin Int(g_{i})$, $\forall i=1,\dots,k$. Then $|Hod(g_{i})|\leq n+1$ and $\max Hod(g_{i})=h(g_{i})$ or $d(g_{i}),\forall i=1,\dots,k$.
By Lemma 3.1, all Weyl modules are Schurian, so condition (1) of Definition 2.1 holds.
To show that each indecomposable projective is good, we verify below $P(h(g_{1}))\in\mathscr{F}$ and the verification of other indecomposable projective modules of hooks being also good modules is similar.
If $|Hod(g_{1})|=n+1$, then $h(g_{1})=d(g_{1})=\max Hod(g_{1})$, $\Delta(h(g_{1}))=P(h(g_{1}))$, so $P(h(g_{1}))\in\mathscr{F}$.
If $|Hod(g_{1})|<n+1$, there are following two cases.
(i) $h(g_{1})=\max Hod(g_{1})$. Then $\Delta(h(g_{1}))=P(h(g_{1}))$, so $P(h(g_{1}))\in\mathscr{F}$.
(ii) $d(g_{1})=\max Hod(g_{1})$. Since $P(h(g_{1}))$ is Schurian, by Lemma 3.2 (3), $P(h(g_{1}))\in\mathscr{F}$.
This finishes the proof. ∎
Theorem 3.3 gives the classification of all quasi-hereditary orderings of simple $A$-modules. We apply Theorem 3.3 to check Example 2.3.
$A$ has only one generator $g=\alpha_{2}\alpha_{1}$ and $Hod(g)=Hod(\alpha_{2}\alpha_{1})=\{1,2,3\}$, $Int(\alpha_{2}\alpha_{1})=\{2\}$. By Theorem 3.3, a simple ordering $\preceq$ is quasi-hereditary if and only if $2$ is not the maximum of $\{1,2,3\}$ under $\preceq$. Therefore, $2\succeq 1\succeq 3$ and $2\succeq 3\succeq 1$ are not in quasi-hereditary, all other 4 simple orderings are quasi-hereditary.
Thanks to Theorem 3.3 and Green-Schroll set $\mathscr{X}$, we can now prove the following criterion to determine whether a Nakayama algebra is quasi-hereditary.
Theorem 3.4.
A Nakayama algebra $A$ is quasi-hereditary if and only if $\mathscr{X}_{A}\neq\emptyset$
Proof.
Necessity. Suppose $\mathscr{X}_{A}=\emptyset$. As $A$ is a quasi-hereditary algebra, $\mathscr{O}(A)\neq\emptyset$. For $\preceq\in\mathscr{O}(A)$, the maximum $s$ under $\preceq$ is not in $\mathscr{X}_{A}$. By Theorem3.3, the elements don’t belong to $\mathscr{X}_{A}$ can’t even be the maximum of a hood, let alone the maximum of $\mathcal{S}(n)$. This leads to contradiction.
Sufficiency. Let $A=KQ_{n}/I$ with $I=<g_{1},\dots,g_{k}>$. Since $\mathscr{X}_{A}\neq\emptyset$, $\ell_{g_{i}}\leq n+1$, $i=1,\dots,k$. Thus $d(g_{i})\notin Int(g_{i})$, $i=1,\dots,k$. Let $x\in\mathscr{X}_{A}$. Consider the following simple ordering $\preceq$”:
$x+1\preceq x+2\preceq\dots\preceq n\preceq 1\preceq\dots\preceq x-1\preceq x$ (1)
According to (1), $\max Int(g_{i})\preceq d(g_{i})$, $i=1,\dots,k$. Namely, $\max Hod(g_{i})\notin Int(g_{i})$, $\forall i=1,\dots,k$. By Theorem 3.3, the simple ordering $\preceq$ defined by (1) is a $q$-ordering. ∎
Theorem 3.4 is the simplest characterization we currently know.
We apply Theorem 3.4 to determine the quasi-heredity of the Nakayama algebra $A$ in Example 2.4. Since $A$ has only one generator $g=\alpha_{2}\alpha_{1}$, it is easy to see that $\mathscr{X}^{1}=\{1,3\}$. Thus $\mathscr{X}\neq\emptyset$. Therefore, $A$ is quasi-hereditary by Theorem 3.4.
3.2. A Brief Survey on Quasi-Heredity Nakayama Algebra
We collect some existing beautiful results for quasi-heredity of Nakayama algebras without proof but with some remarks.
We recall first some notations of Green and Schroll [6]. Let $Q$ be a finite quiver. Let $\mathcal{T}$ be a set of paths of $KQ_{n}$. A vertex $v$ is said to be properly internal to $\mathcal{T}$ if there exist $p,p_{1},p_{2}\in\mathcal{T}$ of length at least 1 such that $p=p_{1}vp_{2}$.
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},...,g_{k}>$. Let $v_{1},\dots,v_{m}$ be some vertices of $Q_{n}$ and $v=v_{1}+...+v_{m}$. The symbol $Q_{\hat{v}}$ denotes the subquiver of $Q_{n}$ deleting all vertices $v_{1},\dots,v_{m}$. The symbol $\mathcal{T}_{\widehat{v_{1}+\dots+v_{m-1}}}$ stands for the set of paths of $KQ_{\hat{v}}$ induced by $g_{i},i=1,...,k$.
Theorem 3.5.
Let $A$ be a Nakayama algebra. The following statements are equivalent:
(1) $A$ is quasi-hereditary.
(2) There is a simple module of projective dimension $2$.
(3) There is a simple ordering of simple modules $v_{1},\dots,v_{n}$ such that for each $i$, the simple module $v_{i}$ is not properly internal to $\mathcal{T}_{\widehat{v_{1}+\dots+v_{i-1}}}$.
(4) $A$ is $S$-connected.
(5) $\mathscr{X}_{A}\neq\emptyset$.
The beautiful condition (2), [3] Proposition $3.1$, can be viewed as the analogy of the well-known fact that algebras of global dimension 2 are quasi-hereditary. Condition (3), [6] Theorem $3.10$, is originally proved to be true for all monomial algebras, so it is a little complicated. Condition (4), [9] Theorem $3.6$, is surprisingly smart, comparing to condition (2). Condition (5), Theorem 3.4 of the present article, is a criterion not involving any algebraic concepts so far.
Using the concepts ”hook” and ”denouement”, we can give a simple characterization of simple modules with projective dimension 1 and 2 over a Nakayama algebra.
Theorem 3.6.
Let $A=KQ_{n}/I$ be a Nakayama algebra and $s\in\mathcal{S}$ non-projective. Then
(1) $pd(s)=1$ if and only if $s$ is not a hook.
(2) $pd(s)=2$ if and only if $s$ is a hook and its denouement belong to $\mathscr{X}_{A}$.
Proof.
(1) $pd(s)=1$ if and only if $radP(s)=P(s+1)$ if and only if $s$ is not a hook.
(2) Necessity. Suppose $pd(s)=2$. By (1), $s$ is the hook $h(g)$ of some generator $g$. Then there is an exact sequence of the following form
[…]
As $s=h(g)$ is simple, $Imf_{2}=Ker\pi=radP(h(g))=(h(g)+1;\ell_{g}-2)$, so $t=h(g)+1$. Since $socP(h(g))=d(g)-1$, $Imf_{1}=Kerf_{2}$, we have $u=d(g)$. Since $f_{1}$ is injective, $P(h(g)+1)$ and $P(d(g))$ have the same socle. Namely, there is no hook in $Int(g)$. Therefore, $d(g)\in\mathscr{X}$.
Sufficiency. Let $s=h(g)$. Since $d(g)\in\mathscr{X}_{A}$, $socP(d(g))=socP(h(g)+1)$. Since $P(d(g))=(d(g);\ell)$, where $\ell$ is the dimension of $P(d(g))$, then $P(h(g)+1)=(h(g)+1;\ell_{g}+\ell-2)$. Since $P(h(g))=(h(g);\ell_{g}-1)$, there is an exact sequence
[…]
where $\pi$ is the canonical surjection with kernel $radP(h(g))$, $\partial_{1}$ is the canonical embedding, $Ker\partial_{0}=Im\partial_{1}$, $Im\partial_{0}=radP(h(g))$. So $pd(h(g))=2$.
Since $pd(s)=0$ if and only if $s$ is a projective module, Theorem 3.6 gains insight into the characteristics of simple modules with projective dimension less than 3. With this understanding, Nakayama algebras with global dimension at least 3 can be characterized via hoods as stated in the following immediate consequence.
Corollary 3.7.
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},\dots,g_{k}>$. Then $gld(A)\geq 3$ if and only if $\bigcup\limits_{i\neq j}[Hod(g_{i})\cap Int(g_{j})]\neq\emptyset$.
Example 3.8.
Let $A=KA_{5}/<\alpha_{2}\alpha_{1},\alpha_{3}\alpha_{2}>$ with
Figure 5. $A_{5}$
By Theorem 3.6, $pd(3)=pd(4)=1$ as $3$, $4$ are not hooks, $pd(2)=2$ as $2=h(\alpha_{3}\alpha_{2})$ and $d(\alpha_{3}\alpha_{2})=4\in\mathscr{X}$. Since $Hod(\alpha_{2}\alpha_{1})\cap Int(\alpha_{3}\alpha_{2})=3$, by Corollary 3.7, $gld(A)\geq 3$. More precisely, since $A$ is quasi-hereditary, by Theorem 3.5 (4), $A$ is S-connected, forcing $pd(1)=3$. Therefore, $gld(A)=3$.
4. applications: Formula of $q(A)$ and $q$-conjecture
In this section, we apply the theory developed in the previous sections to get a general formula for $q(A)$ in the case $A$ is a Nakayama algebra. Based on this formula, the $q$-conjecture is proved to be true for all Nakayama algebras. Unfortunately, we find a counterexample to show that the conjecture fails in general.
4.1. Formula of $q(A)$
So far only one explicit formula for $q(A)$ is given in [5] for non-hereditary algebras. For the convenience of the reader, we restate this formula and involve a short new proof using hoods.
Theorem 4.1.
( [5], Theorem 2) Let $A=KA_{n}/I$ be a directed Nakayama algebra with $I=<g>$ generated by a path $g$, then $q(A)=\frac{2}{\ell_{g}}n!$.
Proof.
Since $|Hod(g)|=\ell_{g}$, we may assume without loss of generality that $Hod(g)=\{i,i+1,\dots,i+\ell_{g}-1\}$. By Theorem 3.3, $i\enspace or\enspace i+\ell_{g}-1=max[i,i+\ell_{g}-1]$. Therefore,
$q(A_{n}/I)=2\tbinom{\ell_{g}}{1}\tbinom{n}{\ell_{g}}(n-\ell_{g})!=\frac{2}{\ell_{g}}n!$
Below we proceed to generalize the formula in Theorem 4.1 to all quasi-hereditary Nakayama algebras.
Let $I=<g_{1},\dots,g_{k}>$ be an admissible ideal $I$ of $KQ_{n}$. Let $\mathcal{T}=\{g_{1},...,g_{k}\}$. The sub-ideal of $I$ missing exactly one generator $g_{i}$ is denoted by $<g_{1},\dots,\hat{g}_{i},\dots,g_{k}>$, namely, $<g_{1},\dots,\hat{g}_{i},\dots,g_{k}>=<\mathcal{T}\setminus\{g_{i}\}>$. Similarly, the sub-ideal $<g_{1},\dots,g_{i-1},g_{i+2},\dots,g_{k}>$ with two generators $g_{i}$ and $g_{i+1}$ deleted is denoted as $<g_{1},\dots,\hat{g}_{i},\hat{g}_{i+1},\dots,g_{k}>$.
Now, we define a new ideal $I_{\hat{x}}$ from $I$ for any simple module $x\in\mathscr{X}$ as follows.
$I_{\hat{x}}=\left\{\begin{array}[]{ll}I,&x\in\mathscr{X}^{0};\\ <g_{1},\dots,\hat{g}_{i},\dots,g_{k}>,&x\in\mathscr{X}^{1}\enspace and\enspace x=h(g_{i})\enspace or\enspace d(g_{i});\\ <g_{1},\dots,\hat{g}_{i},\hat{g}_{i+1},\dots,g_{k}>,&x\in\mathscr{X}^{2}\enspace and\enspace x=d(g_{i})\enspace and\enspace h(g_{i+1})).\end{array}\right.$
Suppose $\preceq$ is a simple ordering of $A$ and $x\in\mathcal{S}$ is a simple $A$-module. We use $\max\limits_{\preceq}\mathcal{S}$ to denote the maximum of $\mathcal{S}$ under $\preceq$. Let
$\mathscr{O}_{x}(A)=\{\preceq|x=\max\limits_{\preceq}\mathcal{S},\preceq\in\mathscr{O}(A)\}$
Denote by $q_{x}(A)=|\mathscr{O}_{x}(A)|$.
Lemma 4.2.
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},\dots,g_{k}>$ and Green-Schroll set $\mathscr{X}_{A}$. Let $x\in\mathcal{S}$. Then
(1) If $x\notin\mathscr{X}_{A}$, $q_{x}(A)=0$.
(2) If $x\in\mathscr{X}_{A}$, $q_{x}(A)=\frac{1}{n}q(KQ_{n}/I_{\hat{x}})$.
Proof.
(1) is a direct consequence of Theorem 3.3.
(2) If $x\in\mathscr{X}_{A}$, there are 3 different cases.
(i) $x\in\mathscr{X}_{A}^{0}$: In this case, $x\notin\bigcup_{i=1}^{k}Hod(g_{i})$ and $I_{\hat{x}}=I$. By Theorem 3.3, $x$ doesn’t influence an ordering $\preceq$ to be quasi-hereditary. Namely, $|\mathscr{O}_{x}(A)|=\frac{1}{n}|\mathscr{O}(A)|$. Therefore, $q_{x}(KQ_{n}/I)=\frac{1}{n}q(KQ_{n}/I)=\frac{1}{n}q(KQ_{n}/I_{\hat{x}})$.
(ii) $x\in\mathscr{X}_{A}^{1}$: This means $x\notin\bigcup_{i=1}^{k}Int(g_{i})$ and there exists $i$ such that $x$ is $h(g_{i})$ or $d(g_{i})$ If $x=\max\mathcal{S}$, then $x=\max Hod(g_{i})$. So the criterion of $Hod(g_{i})$ has been satisfied, $q_{x}(KQ_{n}/I)=q_{x}(KQ_{n}/I_{\hat{x}})$. Since $x\in\mathscr{X}^{0}_{KQ_{n}/I_{\hat{x}}}$ and $(I_{\hat{x}})_{\hat{x}}=I_{\hat{x}}$, using (i), $q_{x}(KQ_{n}/I)=q_{x}(KQ_{n}/I_{\hat{x}})=\frac{1}{n}q(KQ_{n}/I_{\hat{x}})$
(iii) $x\in\mathscr{X}_{A}^{2}$: The proof steps are similar with (ii). ∎
We can now prove the following formula for $q(A)$ via Green-Schroll set.
Theorem 4.3.
Let $A=KQ_{n}/I$ be a Nakayama algebra with Green-Schroll set $\mathscr{X}$. Then
$q(A)=\frac{1}{n}[\sum\limits_{x\in\mathscr{X}^{0}}q(A)+\sum\limits_{x\in\mathscr{X}^{1}}q(KQ_{n}/I_{\hat{x}})+\sum\limits_{x\in\mathscr{X}^{2}}q(KQ_{n}/I_{\hat{x}})]$ (2)
Proof.
By Lemma 4.2,
$q(KQ_{n}/I)$ $=[\sum\limits_{x\in\mathcal{S}}q_{x}(KQ_{n}/I)]$
$=\frac{1}{n}[\sum\limits_{x\in\mathscr{X}}q(KQ_{n}/I_{\hat{x}})]$
$=\frac{1}{n}[\sum\limits_{x\in\mathscr{X}^{0}}q(KQ_{n}/I_{\hat{x}})+\sum\limits_{x\in\mathscr{X}^{1}}q(KQ_{n}/I_{\hat{x}})+\sum\limits_{x\in\mathscr{X}^{2}}q(KQ_{n}/I_{\hat{x}})]$
$=\frac{1}{n}[\sum\limits_{x\in\mathscr{X}^{0}}q(KQ_{n}/I)+\sum\limits_{x\in\mathscr{X}^{1}}q(KQ_{n}/I_{\hat{x}})+\sum\limits_{x\in\mathscr{X}^{2}}q(KQ_{n}/I_{\hat{x}})].$
The following result is a direct consequence of Formula (2).
Corollary 4.4.
Let $A=KQ_{n}/<g>$ be a Nakayama algebra with one generator $g$. Then
$q(A)=\left\{\begin{array}[]{ll}\frac{2}{\ell_{g}}n!,&\ell_{g}\leq n;\\ (n-1)!,&\ell_{g}=n+1;\\ 0,&\ell_{g}>n+1.\end{array}\right.$ (6)
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},\dots,g_{k}>$. The Formula (2) provides an iteration to compute $q(A)$, as explained in the following example.
Example 4.5.
Let $A=KA_{5}/<g_{1},g_{2}>$ with $g_{1}=\alpha_{2}\alpha_{1}$, $g_{2}=\alpha_{3}\alpha_{2}$ be the algebra in Example 3.8. By Theorem 4.3, we know
$q(A)$ $=\frac{1}{5}[q(A)+\sum\limits_{x\in\mathscr{X}^{1}}q(KQ_{n}/I_{\hat{x}})]$
$=\frac{1}{5}[q(A)+q(KQ_{n}/I_{\hat{1}})+q(KQ_{n}/I_{\hat{4}})]$
$=\frac{1}{4}[q(KQ_{n}/I_{\hat{1}})+q(KQ_{n}/I_{\hat{4}})]$
$=\frac{1}{4}[q(KQ_{n}/<g_{2}>)+q(KQ_{n}/<g_{1}>]$
$=\frac{1}{4}(\frac{2}{3}n!+\frac{2}{3}n!)=\frac{1}{3}n!$
The second to last step uses the result of Corollary 4.4.
4.2. Global supremum and $q$-conjecture
In this subsection, we will give the maximum of $q(A)$ for all Nakayama algebras and prove the $q$-conjecture in this case.
Lemma 4.6.
Let $KQ_{n}/I$ be a quasi-hereditary Nakayama Algebra. If $g$ is a path of $KQ_{n}$ and $g\notin I$, then $q(KQ_{n}/I)>q(KQ_{n}/(I+<g>))$.
Proof.
By Theorem 3.3, $\mathscr{O}(KQ_{n}/I)\supseteq\mathscr{O}(KQ_{n}/(I+<g>))$. Since $g\notin I$, there exists an ordering $\preceq\in\mathscr{O}(KQ_{n}/I)$ such that $h(g)\preceq h(g)+1$ and $d(g)\preceq h(g)+1$. Hence, $\max Hod(g)\in Int(g)$ under $\preceq$. This shows $\preceq\notin\mathscr{O}(KQ_{n}/(I+<g>))$. Therefore, $\mathscr{O}(KQ_{n}/I)\supsetneqq\mathscr{O}(KQ_{n}/(I+<g>))$ and $q(KQ_{n}/I)>q(KQ_{n}/(I+<g>))$. ∎
Theorem 4.7.
Let $A=KQ_{n}/I$ be a Nakayama algebra with I an admissible ideal of $KQ_{n}$. Then
$q(A)\leq\frac{2}{3}n!$ (7)
The equality holds if and only if $I$ is a principal ideal generated by a path of length 3. In particular, the q-conjecture is true for all Nakayama algebras.
Proof.
By Lemma 4.6, $q(KQ_{n}/I)>q(KQ_{n}/I+<g>)$, $g\notin I$. Therefore, $q(A)$ reaches its maximum unless $I$ is a principal ideal. By Corollary 4.4, $\max\{q(KQ_{n}/<g>\}=\frac{2}{3}n!$. ∎
Theorem 4.7 shows that the $q$-conjecture holds for all Nakayama algebras. Unfortunately, the following example shows that the conjecture fails in the case that the involved algebra is not monomial.
Example 4.8.
Let $A=KQ_{n}/I$, $I=<\alpha_{2}\alpha_{1}-\alpha_{4}\alpha_{3}>$
[…]
$q(A)=20>\frac{2}{3}\cdot 4!$.