### v2

# Quasi-Hereditary Orderings of Nakayama Algebras

Abstract.
Let $A$ be an algebra with iso-class of simple modules $\mathcal{S}$ of cardinality $n$. A total ordering on $\mathcal{S}$ making every Weyl module Schurian and every indecomposable projective module filtered by the Weyl modules is called to be a quasi-hereditary ordering or $q$-ordering on $A$ and $A$ is a quasi-hereditary algebra under this ordering. The number of $q$-orderings on $A$ is denoted by $q(A)$. To determine whether an ordering on $\mathcal{S}$ is a $q$-ordering is a hard problem. A famous result due to Dlab and Ringel is that $A$ is hereditary if and only if every ordering is a $q$-ordering, equivalently, $q(A)=n!$. The twenty-years old $q$-ordering conjecture claims that $q(A)\leq\dfrac{2}{3}n!$. The present paper proves a very simple criterion for $q$-orderings when $A$ is a Nakayama algebra. This criterion is applied to getting a full classification of all $q$-orderings of $A$ and an explicit iteration formula for $q(A)$, and also a positive proof of the $q$-ordering conjecture for Nakayama algebras.
Supported by National Natural Science Foundation of China under Grant No. 123B1026.
* Corresponding author.
E-mail addresses: zyh@sjtu.edu.cn (Yuehui Zhang), 21-zxq@sjtu.edu.cn (Xiaoqiu Zhong)
Mathematics Subject Classification 2020
16G20, 06A05, 16G10,16D90
Keywords
Quasi-hereditary algebras, Nakayama algebras, q-orderings, Weyl modules
1. Introduction
The theory of quasi-hereditary algebras has been extensively studied since its introduction by E. Cline, B. Parshall, and L. Scott in their seminal paper [1] in 1988. The properties of a quasi-hereditary algebra $A$ are heavily dependent on a specific partial ordering of the iso-classes $\mathcal{S}$ of simple $A$-modules. This ordering ensures that every indecomposable projective module can be filtered by the Weyl modules (or standard modules), which are constructed through a special process associated to the very ordering. Such an ordering is termed a quasi-hereditary ordering — $q$-ordering for short — on $A$. Thus, when referring to a quasi-hereditary algebra, it typically includes specification of the $q$-ordering.
To determine whether an algebra is quasi-hereditary is a very hard problem. The first significant result is given by V. Dlab and C. M. Ringel [2], who showed that global dimension 2 implies quasi-heredity. The integral analogues of this result were obtained by K. Steffen and A. Wiedemann [3] for quasi-hereditary orders. Thus, both hereditary algebras and hereditary orders are quasi-hereditary.
In 1991, C. M. Ringel [4] characterized quasi-hereditary algebras via a special tilting module, the characteristic module. Later, in 1996 and 1997, P. Zhang [5, 6] gave essential improvements to Ringel’s characterization by weakening the characteristic module condition to the existence of a kind of faithful module. He also provided a clever test for the characteristic module by using a property of faithful and partial tilting modules.
Determining which orderings on $\mathcal{S}$ are $q$-orderings for a given quasi-hereditary algebra remains a challenging problem. Dlab and Ringel [7] found that two different partial orderings on $\mathcal{S}$ may yield the same Weyl modules. They defined an equivalence relation on all $q$-orderings and proved that any partial ordering on $\mathcal{S}$ is equivalent to a total ordering. Hence, it suffices to study all total orderings on $\mathcal{S}$. Subsequently, $q$-orderings are generally assumed to be total orderings unless otherwise specified.
This naturally leads to the question: How many $q$-orderings does an arbitrarily given algebra $A$ admit? Denote this number by $q(A)$. Let $|\mathcal{S}|=n$ be the rank of the Grothendieck group of $A$. It is evident that $0\leq q(A)\leq n!$. A classical result, established by Dlab and Ringel, asserts that $A$ is hereditary if and only if $q(A)=n!$, which is equivalent to every ordering on $\mathcal{S}$ being a $q$-ordering [8].
Even two different total orderings may be equivalent. This equivalence relation relates to the concept of ”quasi-hereditary structures” of an algebra $A$, denoted as qh.st($A$), which was defined and systematically studied by M. Flores, Y. Kimura, and B. Rognerud [9]. Precisely, qh.st($A$) is the set of all equivalence classes of all $q$-partial orderings on $\mathcal{S}$. Clearly, qh.st($A$) may be equivalently defined as the set consisting of sets of Weyl modules of all $q$-orderings on $\mathcal{S}$. The notion of quasi-hereditary structures was originally mentioned by K. Coulembier [10]. He proved that if an algebra $A$ admits a simple preserving duality, then $A$ has at most one quasi-hereditary structure.
Quasi-hereditary structures are primarily combinatorial objects, containing limited information about criteria for $q$-orderings and categories of good modules, except for hereditary algebras, even for Nakayama algebras.
The purpose of the present paper is to provide a simple criterion for $q$-orderings for Nakayama algebras. To this end, we need some known facts about Nakayama algebras.
In 1990, M. Uematsu and K. Yamagata demonstrated that Morita equivalence preserves quasi-heredity [11], meaning that if two algebras $A$ and $B$ are Morita equivalent, then $A$ is quasi-hereditary if and only if $B$ is quasi-hereditary. They also established that a non-hereditary Nakayama algebra $A$ is quasi-hereditary if and only if it has a simple module of projective dimension 2.
In 2000, the first author of the present paper and Y. Li proved [12] that $q(A)\leq\frac{2}{3}n!$ for tree-type quasi-hereditary algebras of one generator. Furthermore, they conjectured that this inequality holds for all non-hereditary algebras, namely, they conjectured that $q(A)>\frac{2}{3}n!$ if and only if $q(A)=n!$, which is equivalent to $A$ being hereditary. This conjecture, termed the ”$q$-ordering conjecture”, poses an intriguing challenge in this field. In 2008, the first author, along with L. Wu and C. Gao, proved the $q$-ordering conjecture for $A$ being an $\mathbb{A}$-type algebra with exactly two generators by deriving an explicit formula for $q(A)$ in this context [13].
A remarkable progress is due to E. L. Green and S. Schroll [14]. In 2019, they provided a necessary and sufficient condition for the existence of a $q$-ordering on $\mathcal{S}$ for a monomial algebra $A$ using quiver language. Their theory can be applied to the construction of concrete $q$-orderings by an inductive process on vertices of the underlying quiver of the involved algebra.
Among many approaches on the quasi-heredity of Nakayama algebras, another noteworthy result was established by R. Marczinzik and E. Sen [15] in 2022. They introduced the concept of ”$S$-connected” for a finite-dimensional algebra $A$, defined by the property that the projective dimensions of simple $A$-modules form an interval. They proved that a Nakayama algebra $A$ is $S$-connected if and only if $A$ is quasi-hereditary.
A recent interesting approach due to G. Yuichiro [16] concerned the derived $q$-orderings by permuting the position of simple modules in the given $q$-ordering. He gave a criterion for adjacent transpositions giving quasi-heredity, in terms of homological conditions of Weyl or Verma modules.
The above-mentioned approaches involve limited information about the number of $q$-orderings, even for very simple non-hereditary algebras. Although helpful for deriving explicit formulas for $q(A)$, they are complicated tools for determining whether an ordering of $\mathcal{S}$ is a $q$-ordering, since the computation of all Weyl modules and the checking process of all indecomposable projective modules to be good modules (that is, filtered by Weyl modules, see the definition in the next section) are very demanding tasks.
The key finding of the present paper is a very simple criterion to determine which orderings on $\mathcal{S}$ are $q$-orderings for all Nakayama algebras (Theorem 3.3). We know that usually only a few indecomposable projective modules need to be checked to be good modules. Thus, a much simpler criterion will be obtained if one knows which indecomposable projective modules are always good modules under any ordering of $\mathcal{S}$, so only the others need to be considered. Fortunately, we find two kinds of simple modules called ”hooks” and ”denouements” that provide such a criterion for all Nakayama algebras. This surprisingly simple criterion can be used to obtain the complete classification of all $q$-orderings, thus can be further applied to classify quasi-hereditary structures. Other applications of hooks and denouements include characterizations of simple modules with projective dimensions 1 or 2 over Nakayama algebras, leading to a characterization of Nakayama algebras with global dimensions greater than 2 (Proposition 2.2).
The other key finding of the present paper is an easy characterization of the quasi-heredity of a Nakayama algebra $A$. We find that the quasi-heredity of $A$ is determined by a subset $\mathscr{X}_{A}$, the $Q$-set of $A$, of some special simple modules. We show that $A$ is quasi-hereditary if and only if its $Q$-set is nonempty (Theorem 3.4). The $Q$-set is so simple that it can be found quickly for any Nakayama algebra, even without any calculations. We also provide an iteration formula for $q(A)$ by means of the $Q$-set. The criterion is further applied to computing the global supremum of $q(A)$ for all Nakayama algebras, thereby affirming the $q$-ordering conjecture for Nakayama algebras (Theorem 4.7). Finally, we include a counterexample to demonstrate that the $q$-ordering conjecture may fail when $A$ is not monomial.
Unless otherwise specified, all algebras discussed in this paper are assumed to be (connected) finite-dimensional over a fixed algebraically closed field $K$, with the unit element of an algebra $A$ denoted as $1_{A}$. For any set $X$, its cardinality is denoted by $|X|$.
2. Preliminaries
2.1. Quasi-hereditary Algebra
Let $A$ be an algebra, and let $\Lambda$ be a finite poset in bijective correspondence with the set of iso-classes $\mathcal{S}$ of simple $A$-modules. The poset $\Lambda$ is referred to as the weight poset of $A$, with its elements being termed weights of $A$. For each weight $\lambda\in\Lambda$, let $S_{\lambda}$ denote the corresponding simple module and $P(\lambda)$ denote its projective cover. Define $\Delta(\lambda)$ as the maximal factor module of $P(\lambda)$ with composition factors of the form $S_{\mu}$, where $\mu\preceq\lambda$. Denote by $\Delta$ the full subcategory of $A$-mod consisting of all $\Delta(\lambda)$, where $\lambda\in\Lambda$. The modules in $\Delta$ are termed Weyl modules.
Let ${\mathscr{F}}(\Delta)$ denote the class of all $A$-modules filtered by $\Delta$, meaning there exists a filtration
$0=M_{t}\subset M_{t-1}\subset\cdots\subset M_{1}\subset M_{0}=M$
such that each factor module $M_{i-1}/M_{i}$ is isomorphic to some Weyl module, for all $1\leq i\leq t$. Modules in ${\mathscr{F}}(\Delta)$ are referred to as good modules. For simplicity, ${\mathscr{F}}(\Delta)$ is denoted simply as ${\mathscr{F}}$.
Definition 2.1.
The algebra $A$ is said to be quasi-hereditary with respect to $(\Lambda,\preceq)$ if for each $\lambda\in\Lambda$ we have:
  1. (1)
    $End_{A}(\Delta(\lambda))\simeq K$;
  2. (2)
    $P(\lambda)\in{\mathscr{F}}$.
By $(A,\Lambda)$ we denote a quasi-hereditary algebra $A$ with weight poset $\Lambda$. We often omit the weight poset $\Lambda$ if it is obvious in the context.
A module $M$ satisfying condition (1) of 2.1 (i.e. $End_{A}M\simeq K$) is usually called to be Schurian, so an algebra $A$ is quasi-hereditary if and only if all Weyl modules are Schurian and all projective modules are good modules.
2.2. Nakayama Algebras
According to [11], Proposition 2.3, Morita equivalent algebras share the same quasi-heredity properties. Additionally, it is well-known that any finite-dimensional algebra is Morita equivalent to a quotient algebra of a path algebra defined by a finite quiver. Hence, we will concentrate our investigation on the quasi-heredity of Nakayama algebras by categorizing them into directed and cyclic types. Directed Nakayama algebras are those Morita equivalent to algebras defined by directed quivers, while cyclic Nakayama algebras are those Morita equivalent to algebras defined by cyclic quivers.
2.2.1. Directed Nakayama Algebras
For any $n\geq 2$, $A_{n}$ denotes the following quiver:
Figure 1. $A_{n}$
The arrow from vertex $j$ to $j+1$ is denoted by $\alpha_{j}$, where $11\leq j\leq n-$ 1.
The oriented quiver $A_{n}$ is commonly referred to as the linearly ordered Dynkin diagram with $n$ vertices.
An algebra $A$ is termed a directed Nakayama algebra of $A_{n}$ type if $A$ is a factor algebra $KA_{n}/I$ of the quiver algebra $KA_{n}$, where $I$ is an admissible and two-sided ideal of $KA_{n}$.
2.2.2. Cyclic Nakayama Algebra
For any $n\geq 2$, $\widetilde{A}_{n}$ denotes the following quiver:
Figure 2. $\widetilde{A}_{n}$
In the cyclic case, the arrows from vertex $j$ to $j+1$ and from vertex $n$ to $1$ are denoted by $\alpha_{j}$, where $11\leq j\leq n-$ 1, and $\alpha_{n}$ respectively.
The oriented quiver $\widetilde{A}_{n}$ is commonly referred to as the Euclidean diagram with $n$ vertices.
An algebra $A$ is considered a cyclic Nakayama algebra if $A$ is a factor algebra $K\widetilde{A}_{n}/I$ of the path algebra $K\widetilde{A}_{n}$ with $I$ being an admissible and two-sided ideal of $K\widetilde{A}_{n}$.
Cyclic Nakayama algebra is also known as Euclidean type Nakayama algebra.
2.3. Notations of Simple Modules, Hooks and Denouements of Paths
Let $Q$ be any finite quiver, and let $A=KQ/I$ be a finite-dimensional algebra with $I$ being an admissible ideal of $KQ$. For simplicity, we denote the simple $A$-module $S_{x}$ corresponding to the vertex $x$ (i.e., $S_{x}=A/AeA$ with $e=1_{A}-x$) simply as $x$. Thus, the symbol $x$ simultaneously represents a number, a vertex, and a simple module. Consequently, the weight poset $\Lambda$ of $A$ can be chosen as the set $\mathcal{S}$ of all simple $A$-modules. This notation system prevents confusion when we use $\preceq$ (or its reverse order $\succeq$) to denote the partial order of the weight poset $\Lambda$.
Consider $g\in I$ to be a path of length(= the number of arrows $+1$) $\ell_{g}\geq 3$ in $KQ$. The origin vertex of $g$ is referred to as the ”hook” of $g$, denoted as $h(g)$, while the end vertex of $g$ is termed the ”denouement” of $g$, denoted as $d(g)$. The ordered multiple set $\{h(g),h(g)+1,\dots,d(g)\}$ of simple $A$-modules associated with $g$ is called the ”hood” of $g$ and denoted as $Hod(g)$. The multiple set $Hod(g)\setminus\{h(g),d(g)\}$ of simple $A$-modules is denoted by $Int(g)$, where the elements of $Int(g)$ are referred to as interior simple modules of $g$. It is noteworthy that $|Hod(g)|=\ell_{g}$ and $|Int(g)|=\ell_{g}-2$.
Given any simple ordering $\preceq$ of $\mathcal{S}$. Denote by $\max Hod(g)$ the maximum of $Hod(g)$ with respect to $\preceq$.
There are several characterizations of simple modules with projective dimensions $1$ and $2$. For instance, R. Marczinzik, M. Rubey, and C. Stump [17] utilize Kupisch series to achieve such characterizations. By employing the concepts of “hooks” and “denouements,” we can also provide a straightforward characterization of simple modules with projective dimensions $1$ and $2$ over a Nakayama algebra. This, in turn, leads to a characterization of Nakayama algebras with global dimension greater than $2$.
Proposition 2.2.
Let $A=KQ_{n}/I$ be a Nakayama algebra and $s\in\mathcal{S}$ non-projective. Then
  1. (1)
    $pd(s)=1$ if and only if $s$ is not a hook.
  2. (2)
    $pd(s)=2$ if and only if $s$ is a hook and its denouement belong to $\mathscr{X}_{A}$.
  3. (3)
    Let $I=<g_{1},\dots,g_{k}>$. Then $gld(A)\geq 3$ if and only if $\bigcup\limits_{i\neq j}[Hod(g_{i})\cap Int(g_{j})]\neq\emptyset$.
Proof.
  1. (1)
    $pd(s)=1$ if and only if $radP(s)=P(s+1)$ if and only if $s$ is not a hook.
  2. (2)
    1. (a)
      Necessity. Suppose $pd(s)=2$. By (1), $s$ is the hook $h(g)$ of some generator $g$. Then there is an exact sequence of the following form
[…]
      As $s=h(g)$ is simple, $Imf_{2}=Ker\pi=radP(h(g))=(h(g)+1;\ell_{g}-2)$, so $t=h(g)+1$. Since $socP(h(g))=d(g)-1$, $Imf_{1}=Kerf_{2}$, we have $u=d(g)$. Since $f_{1}$ is injective, $P(h(g)+1)$ and $P(d(g))$ have the same socle. Namely, there is no hook in $Int(g)$. Therefore, $d(g)\in\mathscr{X}$.
    2. (b)
      Sufficiency. Let $s=h(g)$. Since $d(g)\in\mathscr{X}_{A}$, $socP(d(g))=socP(h(g)+1)$. Since $P(d(g))=(d(g);\ell)$, where $\ell$ is the dimension of $P(d(g))$, then $P(h(g)+1)=(h(g)+1;\ell_{g}+\ell-2)$. Since $P(h(g))=(h(g);\ell_{g}-1)$, there is an exact sequence
[…]
      where $\pi$ is the canonical surjection with kernel $radP(h(g))$, $\partial_{1}$ is the canonical embedding, $Ker\partial_{0}=Im\partial_{1}$, $Im\partial_{0}=radP(h(g))$. So $pd(h(g))=2$.
  3. (3)
    It is well known that the global dimension of artinian algebra is the maximum of projective dimensions of all simple modules, the conclusion follows from (1) and (2), and $pd(s)=0$ if and only if $s$ is projective.
According to [11], a non-hereditary Nakayama algebra $A$ is quasi-hereditary if and only if it has a simple module of projective dimension $2$. Now, one can use Proposition 2.2 to quickly identify all simple modules of projective dimension $2$.
Example 2.3.
Let $A=KA_{5}/<\alpha_{2}\alpha_{1},\alpha_{3}\alpha_{2}>$ with
Figure 3. $A_{5}$
By Proposition 2.2, $pd(3)=pd(4)=1$ as $3$, $4$ are not hooks, $pd(2)=2$ as $2=h(\alpha_{3}\alpha_{2})$ and $d(\alpha_{3}\alpha_{2})=4\in\mathscr{X}$. Since $Hod(\alpha_{2}\alpha_{1})\cap Int(\alpha_{3}\alpha_{2})=3$, by Proposition 2.2, $gld(A)\geq 3$. More precisely, since $A$ is quasi-hereditary, $A$ is S-connected [15], forcing $pd(1)=3$. Therefore, $gld(A)=3$.
2.4. Definition of $Q$-set for Nakayama Algebras
We will use $Q_{n}$ to denote both the quiver $A_{n}$ and $\widetilde{A}_{n}$. Thus, an arbitrarily given Nakayama algebra is Morita equivalent to $KQ_{n}/I$ for some $n$ with an admissible ideal $I$.
Since Morita equivalence preserves quasi-heredity, to analyze $q$-orderings of Nakayama algebras, it suffices to focus on Nakayama algebras of the form $KQ_{n}/I$. Each generator $g=\alpha_{d(g)-1}\alpha_{d(g)-2}\cdots\alpha_{h(g)}$ of $I$ is a path from its origin vertex $h(g)$ to its end vertex $d(g)$ of length $\ell_{g}\geq 3$, where $d(g)=h(g)+\ell_{g}-1$. In the event that a number $v$ referring to a vertex or a simple module exceeds $n$, we understand it as the vertex or simple module $v\,\text{mod}(n)$.
Note that any admissible ideal $I$ of $KQ_{n}$ is generated by a minimal set $\{g_{1},\dots,g_{k}\}$ of generators, where each $g_{i}$ is a path from its origin vertex $v_{i}$ and $v_{1}<v_{2}<\dots<v_{k}$. Throughout, when we write an admissible ideal as $I=\{g_{1},\dots,g_{k}\}$, the set $\{g_{1},\dots,g_{k}\}$ is always supposed to be minimal.
Definition 2.4.
For a Nakayama algebra $A=KQ_{n}/I$ with $I=\langle g_{1},\dots,g_{k}\rangle$, $\mathscr{X}_{A}$ or simply $\mathscr{X}$, is the set $\mathcal{S}-\bigcup_{i=1}^{k}\text{Int}(g_{i})$. Let $\mathscr{X}^{0}$, $\mathscr{X}^{1}$, $\mathscr{X}^{2}$, respectively, be the subset of $\mathscr{X}$ whose elements do not belong to any hood, are either hooks or denouements of some hoods but not both, and are both hooks and denouements of some hoods, respectively.
Note that $\mathscr{X}^{0}$, $\mathscr{X}^{1}$, $\mathscr{X}^{2}$ form a partition of $\mathscr{X}$, namely, they are disjoint subsets of $\mathscr{X}$ and $\mathscr{X}=\mathscr{X}^{0}\cup\mathscr{X}^{1}\cup\mathscr{X}^{2}$.
As we will see in Section 3 (Theorem 3.4), the quasi-heredity of a Nakayama algebra $A$ is completely determined by $\mathscr{X}_{A}$. The computation of $q(A)$ is also highly dependent on $\mathscr{X}_{A}$.
2.5. Notations for indecomposable modules over Nakayama Algebras
Since Nakayama algebras are serial, each indecomposable module possesses precisely one composition series (from its top to its socle), completely determined by its top and dimension (length). To denote the unique indecomposable projective module with top $t$ and dimension $\ell$, we employ the symbol $(t;\ell)$. Hence, the indecomposable projective module $P(j)$, where $top(P(j))=j$, is exactly the module $(j;\ell)$, where $\ell=\dim_{K}P(j)$. Furthermore, from this representation, it is evident that $radP(j)=(j+1;\ell-1)$.
To deal with the $q$-orderings, note that the condition (1) in the Definition 2.1 is automatically satisfied for all directed algebras; for a cyclic Nakayama algebra $KQ_{n}/I$, to assure the condition (1), the indecomposable projective module covering the largest simple module (according to the corresponding $q$-ordering) must be multiplicity free, thus its length has to be at most $n$, forcing the shortest length of all generators of $I$ is at most $n$.
Another easy fact of a directed Nakayama algebra is that it has a unique simple projective module $P(n)$ and obviously $\Delta(n)=P(n)$ under all simple orderings. So, $P(n)$ always satisfies the condition (2) in the Definition 2.1.
Example 2.5.
Let $A=KA_{3}/<\alpha_{2}\alpha_{1}>$ with
Figure 4. $A_{3}$
Choose a simple ordering on the weight poset $\Lambda=\mathcal{S}$ as $1\succeq 2\succeq 3$, then $P(k)=\Delta(k)$, for $k=1,2,3$, so $1\succeq 2\succeq 3$ is a $q$-ordering. Similarly, one checks easily that $1\succeq 3\succeq 2$ is also a $q$-ordering. However, if one chooses the weight poset $\Lambda$ with ordering the simple modules as $2\succeq 1\succeq 3$, then $P(k)=\Delta(k)$, for $k=2,3$, but $\Delta(k)=1$ and $P(1)\not\in\mathscr{F}$, so $2\succeq 1\succeq 3$ is not a $q$-orderings. In fact, check other 3 orderings, $3\succeq 1\succeq 2$ and $3\succeq 2\succeq 1$ are $q$-ordering, $2\succeq 3\succeq 1$ is not. So $q(A)=4$.
Example 2.6.
Let $A=K\tilde{A}_{3}/<\alpha_{2}\alpha_{1}>$ with
Figure 5. $\tilde{A}_{3}$
Similar to Example 2.5, one can easily get $q(A)=4$. So q-conjecture holds in this case.
3. Quasi-hereditary Nakayama Algebras
In this section, we establish a criterion for $q$-orderings over Nakayama algebras through the use of hoods. Subsequently, we derive a concise characterization for the quasi-heredity of Nakayama algebras, where the non-emptiness of the $Q$-set serves as a necessary and sufficient condition.
3.1. Quasi-hereditary orderings of Nakayama Algebras
Let $g$ be a path of $KQ$. If the cardinality $|Hod(g)|$ of the hood $Hod(g)$ is smaller enough, then the indecomposable projective module $P(h(g))$ is Schurian, forcing the Weyl module $\Delta(h(g))$ is also Schurian, as stated in the following lemma.
Lemma 3.1.
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},\dots,g_{k}>$. If $|Hod(g_{i})|\leq n+1$, $i=1,\dots,k$, then every Weyl module $\Delta(h(g_{i}))$ is Schurian under any simple ordering.
Hooks and denouements play key roles in the following result.
Lemma 3.2.
Let $A=KQ_{n}/I$ be a Nakayama algebra, $1\leq k\leq n$. Given any simple ordering $\preceq$ with good module category $\mathscr{F}$.
(1) If $radP(k)\in\mathscr{F}$, then $P(k)\in\mathscr{F}$.
(2) If k is not a hook and $P(k+1)\in\mathscr{F}$, then $P(k)\in\mathscr{F}$.
(3) Let g be a generator of $I$ with hook $h(g)=k$ and denouement $d(g)$. If $d(g)=\max Hod(g)$ and $P(k)$ is Schurian, then $P(k)\in\mathscr{F}$.
Proof.
(1) If $P(k)=(k;\ell)$, then $radP(k)=(k+1;\ell-1)$. Suppose $\Delta(k)=(k;\ell^{\prime})$, $1\leq\ell^{\prime}\leq\ell$. There are 3 cases.
(i) $\ell^{\prime}=\ell$.
Then $\Delta(k)=(k;\ell)=P(k)$, thus $P(k)\in\mathscr{F}$.
(ii) $\ell^{\prime}=1$.
Then $\Delta(k)=k$. It is obvious that $P(k)\in\mathscr{F}$ since $P(K)$ is filtered by good modules $radP(k)$ and $\Delta(k)$.
(iii) $1<\ell^{\prime}<\ell$.
Thus $k=\max[k,k+\ell^{\prime}-1]$. So, $k+\ell^{\prime}\not\in\Delta(i),\forall\enspace k\leq i\leq k+\ell^{\prime}-1$ Since $radP(k+1)\in\mathscr{F}$, there exists a module $M=(k+\ell^{\prime};\ell-\ell^{\prime})\in\mathscr{F}$. Thus $P(k)\in\mathscr{F}$ as $P(k)$ is filtered by good modules $M$ and $\Delta(k)$.
(2) If $k$ is a hook, then $radP(k)=P(k+1)$. So $radP(k)\in\mathscr{F}$. By (1), $P(k)\in\mathscr{F}$.
(3) Let $d(g)=\max Hod(g)$ and $P(k)$ Schurian. Then $P(k)=(k;\ell_{g}-1)$. Since $d(g)=\max Hod(g)$, each Weyl module $\Delta(i)$ with $k\leq i\leq\ell_{g}-1$ is completely determined by the restriction of $\preceq$ on $Hod(g)$ and irrelevant to any simple modules outside $Hod(g)\setminus\{d(g)\}$. Now the full subcategory $\mathcal{C}$ of $modA$ generated by the simple modules of $Hod(g)\setminus\{d(g)\}$ is Morita equivalent to (indeed isomorphic to) the module category of the hereditary algebra $KA_{\ell_{g}-1}$, thus $\mathcal{C}$ is also a hereditary abelian category. Therefore every simple ordering on $Hod(g)\setminus\{d(g)\}$ is a $q$-ordering, forcing $P(k)\in\mathscr{F}$. ∎
Now we can prove the following criterion for $q$-orderings.
Theorem 3.3.
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},\dots,g_{k}>$. Let $\preceq$ be a simple ordering of $A$. Then $\preceq$ is quasi-hereditary if and only if $\max Hod(g_{i})\notin Int(g_{i})$, $\forall i=1,\dots,k$.
Proof.
Sufficiency. Assume there exists a $Hod(g)$ whose maximum $o\in Int(g)$ under $\preceq$. There are following two cases.
(1) If $o=h(g)$ or $o=d(g)$. Then the multiplicity of the simple module $o$ in $\Delta(o)$ is at least 2, so $dim_{K}End_{A}(\Delta(o))\geq 2$, so the condition (1) in Definition 2.1 of quasi-hereditary algebra fails, contradicting that $\preceq$ is a $q$-ordering.
(2) If $o\neq h(g),d(g)$. $\forall i\in Hod(g)$, $o\notin\Delta(i)$ except $\Delta(o)$. By direct computation, there exists $j\geq 0$, $soc\Delta(o)=d(g)+j,socP(h(g))=d(g)-1$. Note that $o\succeq h(g),o\in P(h(g))$, so $P(h(g))$ does not admit a $\Delta$-filtration, thus $P(h(g))\notin\mathscr{F}$, so the condition (2) in Definition 2.1 fails, which contradicts that $\preceq$ is a $q$-ordering.
Necessity. Assume $\preceq$ satisfies $\max Hod(g_{i})\notin Int(g_{i})$, $\forall i=1,\dots,k$. Then $|Hod(g_{i})|\leq n+1$ and $\max Hod(g_{i})=h(g_{i})$ or $d(g_{i}),\forall i=1,\dots,k$.
By Lemma 3.1, if $s$ is a hook, then $\Delta(s)$ is Schurian. Since $\max Hod(g_{i})=h(g_{i})$ or $d(g_{i}),\forall i=1,\dots,k$, $\Delta(s)$ is Schurian for all $s\in Int(g_{i})$. Above all, condition (1) of Definition 2.1 holds.
To show that each indecomposable projective is good, we verify below $P(h(g_{1}))\in\mathscr{F}$ and the verification of other indecomposable projective modules of hooks being also good modules is similar.
If $|Hod(g_{1})|=n+1$, then $h(g_{1})=d(g_{1})=\max Hod(g_{1})$, $\Delta(h(g_{1}))=P(h(g_{1}))$, so $P(h(g_{1}))\in\mathscr{F}$.
If $|Hod(g_{1})|<n+1$, there are following two cases.
(i) $h(g_{1})=\max Hod(g_{1})$. Then $\Delta(h(g_{1}))=P(h(g_{1}))$, so $P(h(g_{1}))\in\mathscr{F}$.
(ii) $d(g_{1})=\max Hod(g_{1})$. Since $P(h(g_{1}))$ is Schurian, by Lemma 3.2 (3), $P(h(g_{1}))\in\mathscr{F}$.
So all $P(h(g_{i}))\in\mathscr{F}$.
Now the other indecomposable projective modules being good modules follows from lemma 3.2 (2).
This finishes the proof. ∎
Theorem 3.3 provides a very simple criterion to determine which orderings on $\mathcal{S}$ are $q$-orderings for all Nakayama algebras. Specifically, to verify whether a simple ordering of a Nakayama algebra $A$ is a $q$-ordering, it suffices to check the restriction of this ordering on each hood of $A$. The restriction on each hood requires that its maximal simple module does not belong to its interior. Consequently, the verification of $q$-orderings is reduced to examining the corresponding simple orderings of one-generator Nakayama algebras.
We apply Theorem 3.3 to check Example 2.5.
$A$ has only one generator $g=\alpha_{2}\alpha_{1}$ and $Hod(g)=Hod(\alpha_{2}\alpha_{1})=\{1,2,3\}$, $Int(\alpha_{2}\alpha_{1})=\{2\}$. By Theorem 3.3, a simple ordering $\preceq$ is quasi-hereditary if and only if $2$ is not the maximum of $\{1,2,3\}$ under $\preceq$. Therefore, $2\succeq 1\succeq 3$ and $2\succeq 3\succeq 1$ are not in quasi-hereditary, all other 4 simple orderings are quasi-hereditary.
Thanks to Theorem 3.3 and $Q$-set, we can now prove the following criterion to determine whether a Nakayama algebra is quasi-hereditary.
Theorem 3.4.
A Nakayama algebra $A$ is quasi-hereditary if and only if $\mathscr{X}_{A}\neq\emptyset$
Proof.
Necessity. Suppose $\mathscr{X}_{A}=\emptyset$. As $A$ is a quasi-hereditary algebra, $\mathscr{O}(A)\neq\emptyset$. For $\preceq\in\mathscr{O}(A)$, the maximum $s$ under $\preceq$ is not in $\mathscr{X}_{A}$. By Theorem 3.3, the elements don’t belong to $\mathscr{X}_{A}$ can’t even be the maximum of a hood, let alone the maximum of $\mathcal{S}(n)$. This leads to contradiction.
Sufficiency. Let $A=KQ_{n}/I$ with $I=<g_{1},\dots,g_{k}>$. Since $\mathscr{X}_{A}\neq\emptyset$, $\ell_{g_{i}}\leq n+1$, $i=1,\dots,k$. Thus $d(g_{i})\notin Int(g_{i})$, $i=1,\dots,k$. Let $x\in\mathscr{X}_{A}$. Consider the following simple ordering $\preceq$”:
$x+1\preceq x+2\preceq\dots\preceq n\preceq 1\preceq\dots\preceq x-1\preceq x$ (1)
According to (1), $\max Int(g_{i})\preceq d(g_{i})$, $i=1,\dots,k$. Namely, $\max Hod(g_{i})\notin Int(g_{i})$, $\forall i=1,\dots,k$. By Theorem 3.3, the simple ordering $\preceq$ defined by (1) is a $q$-ordering. ∎
Theorem 3.4 seems to be the simplest characterization we currently know, since $Q$-set is so easy to be found for any Nakayama algebras even without any calculations.
We apply Theorem 3.4 to determine the quasi-heredity of the Nakayama algebra $A$ in Example 2.6. Since $A$ has only one generator $g=\alpha_{2}\alpha_{1}$, it is easy to see that $\mathscr{X}^{1}=\{1,3\}$. Thus $\mathscr{X}\neq\emptyset$. Therefore, $A$ is quasi-hereditary by Theorem 3.4.
3.2. A Brief Survey on Quasi-Hereditary Nakayama Algebra
We include a very brief survey in this subsection of quasi-hereditary Nakayama algebras without proof but with some remarks.
We recall first some notations of Green and Schroll [14]. Let $Q_{n}$ be a finite quiver. Let $\mathcal{T}$ be a set of paths of $KQ_{n}$. A vertex $v$ is said to be properly internal to $\mathcal{T}$ if there exist $p,p_{1},p_{2}\in\mathcal{T}$ of length at least 1 such that $p=p_{1}vp_{2}$.
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},...,g_{k}>$. Let $v_{1},\dots,v_{m}$ be some vertices of $Q_{n}$ and $v=v_{1}+...+v_{m}$. The symbol $Q_{\hat{v}}$ denotes the subquiver of $Q_{n}$ deleting all vertices $v_{1},\dots,v_{m}$. The symbol $\mathcal{T}_{\widehat{v_{1}+\dots+v_{m-1}}}$ stands for the set of paths of $KQ_{\hat{v}}$ induced by $g_{i},i=1,...,k$.
The following theorem is a collection of existing beautiful results for quasi-heredity of Nakayama algebras.
Theorem 3.5.
Let $A$ be a Nakayama algebra. The following statements are equivalent:
(1) $A$ is quasi-hereditary.
(2) There is a simple module of projective dimension $2$, or $A$ is hereditary.
(3) There is a simple ordering of simple modules $v_{1},\dots,v_{n}$ such that for each $i$, the simple module $v_{i}$ is not properly internal to $\mathcal{T}_{\widehat{v_{1}+\dots+v_{i-1}}}$.
(4) $A$ is $S$-connected.
(5) $\mathscr{X}_{A}\neq\emptyset$.
Condition (2) ([11] Proposition $3.1$) can be viewed as the analogy of the well known fact that algebras of global dimension 2 are quasi-hereditary. It is so beautiful that one may forget its difficulty for the computation of the projective dimensions of simple modules, which is equivalent to the computation of the global dimension of $A$. Though condition (3) ([14] Theorem $3.10$) is complicated both theoretically and technically, it is very useful since it is originally proved for all monomial algebras. Condition (4) ([15] Theorem $3.6$) is as beautiful as condition (2), with application difficulty also as condition (2), since it involves the computation of the projective dimensions of all simple modules. Condition (5), Theorem 3.4 of the present paper, provides a criterion that does not involve any algebraic concepts. It is so straightforward to apply that one only needs to examine the starting points of all generators of $A$ to determine the quasi-heredity of $A$, without requiring any computation.
4. applications: Formula of $q(A)$ and $q$-conjecture
In this section, we give two applications of the theory developed in the previous sections. One is a general formula for $q(A)$ in the case $A$ is a Nakayama algebra. The other is a proof to the $q$-conjecture for all Nakayama algebras. We also include a counterexample to show that the conjecture may fail in case that algebras involved are not monomial.
4.1. Formula of $q(A)$
So far only one explicit formula for $q(A)$ is given in [13] for non-hereditary algebras. For the convenience of the reader, we restate this formula and involve a short new proof using hoods.
Theorem 4.1.
( [13], Theorem 2) Let $A=KA_{n}/I$ be a directed Nakayama algebra with $I=<g>$ generated by a path $g$, then $q(A)=\frac{2}{\ell_{g}}n!$.
Proof.
Since $|Hod(g)|=\ell_{g}$, we may assume without loss of generality that $Hod(g)=\{i,i+1,\dots,i+\ell_{g}-1\}$. By Theorem 3.3, $i\enspace or\enspace i+\ell_{g}-1=max[i,i+\ell_{g}-1]$. Therefore,
$q(A_{n}/I)=2\tbinom{\ell_{g}}{1}\tbinom{n}{\ell_{g}}(n-\ell_{g})!=\frac{2}{\ell_{g}}n!$
Below we proceed to generalize the formula in Theorem 4.1 to all quasi-hereditary Nakayama algebras.
Let $I=<g_{1},\dots,g_{k}>$ be an admissible ideal $I$ of $KQ_{n}$. Let $\mathcal{T}=\{g_{1},...,g_{k}\}$. The sub-ideal of $I$ missing exactly one generator $g_{i}$ is denoted by $<g_{1},\dots,\hat{g}_{i},\dots,g_{k}>$, namely, $<g_{1},\dots,\hat{g}_{i},\dots,g_{k}>=<\mathcal{T}\setminus\{g_{i}\}>$. Similarly, the sub-ideal $<g_{1},\dots,g_{i-1},g_{i+2},\dots,g_{k}>$ with two generators $g_{i}$ and $g_{i+1}$ deleted is denoted as $<g_{1},\dots,\hat{g}_{i},\hat{g}_{i+1},\dots,g_{k}>$.
Now, we define a new ideal $I_{\hat{x}}$ from $I$ for any simple module $x\in\mathscr{X}$ as follows.
$I_{\hat{x}}=\left\{\begin{array}[]{ll}I,&x\in\mathscr{X}^{0};\\ <g_{1},\dots,\hat{g}_{i},\dots,g_{k}>,&x\in\mathscr{X}^{1}\enspace and\enspace x=h(g_{i})\enspace or\enspace d(g_{i});\\ <g_{1},\dots,\hat{g}_{i},\hat{g}_{i+1},\dots,g_{k}>,&x\in\mathscr{X}^{2}\enspace and\enspace x=d(g_{i})\enspace and\enspace h(g_{i+1})).\end{array}\right.$
Suppose $\preceq$ is a simple ordering of $A$ and $x\in\mathcal{S}$ is a simple $A$-module. We use $\max\limits_{\preceq}\mathcal{S}$ to denote the maximum of $\mathcal{S}$ under $\preceq$. Let
$\mathscr{O}_{x}(A)=\{\preceq|x=\max\limits_{\preceq}\mathcal{S},\preceq\in\mathscr{O}(A)\}$
Denote by $q_{x}(A)=|\mathscr{O}_{x}(A)|$.
Lemma 4.2.
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},\dots,g_{k}>$. Let $x\in\mathcal{S}$. Then
(1) If $x\notin\mathscr{X}_{A}$, $q_{x}(A)=0$.
(2) If $x\in\mathscr{X}_{A}$, $q_{x}(A)=\frac{1}{n}q(KQ_{n}/I_{\hat{x}})$.
Proof.
(1) is a direct consequence of Theorem 3.3.
(2) If $x\in\mathscr{X}_{A}$, there are 3 different cases.
(i) $x\in\mathscr{X}_{A}^{0}$: In this case, $x\notin\bigcup_{i=1}^{k}Hod(g_{i})$ and $I_{\hat{x}}=I$. By Theorem 3.3, $x$ doesn’t influence an ordering $\preceq$ to be quasi-hereditary. Namely, $|\mathscr{O}_{x}(A)|=\frac{1}{n}|\mathscr{O}(A)|$. Therefore, $q_{x}(KQ_{n}/I)=\frac{1}{n}q(KQ_{n}/I)=\frac{1}{n}q(KQ_{n}/I_{\hat{x}})$.
(ii) $x\in\mathscr{X}_{A}^{1}$: This means $x\notin\bigcup_{i=1}^{k}Int(g_{i})$ and there exists $i$ such that $x$ is $h(g_{i})$ or $d(g_{i})$ If $x=\max\mathcal{S}$, then $x=\max Hod(g_{i})$. So the criterion of $Hod(g_{i})$ has been satisfied, $q_{x}(KQ_{n}/I)=q_{x}(KQ_{n}/I_{\hat{x}})$. Since $x\in\mathscr{X}^{0}_{KQ_{n}/I_{\hat{x}}}$ and $(I_{\hat{x}})_{\hat{x}}=I_{\hat{x}}$, using (i), $q_{x}(KQ_{n}/I)=q_{x}(KQ_{n}/I_{\hat{x}})=\frac{1}{n}q(KQ_{n}/I_{\hat{x}})$
(iii) $x\in\mathscr{X}_{A}^{2}$: The proof steps are similar with (ii). ∎
We can now prove the following formula for $q(A)$ via $\mathscr{X}_{A}$.
Theorem 4.3.
Let $A=KQ_{n}/I$ be a Nakayama algebra and $\mathscr{X}_{A}$ be abbreviated as $\mathscr{X}$. Then
$q(A)=\frac{1}{n}[\sum\limits_{x\in\mathscr{X}^{0}}q(A)+\sum\limits_{x\in\mathscr{X}^{1}}q(KQ_{n}/I_{\hat{x}})+\sum\limits_{x\in\mathscr{X}^{2}}q(KQ_{n}/I_{\hat{x}})]$ (2)
Proof.
By Lemma 4.2,
$q(KQ_{n}/I)$ $=[\sum\limits_{x\in\mathcal{S}}q_{x}(KQ_{n}/I)]$
$=\frac{1}{n}[\sum\limits_{x\in\mathscr{X}}q(KQ_{n}/I_{\hat{x}})]$
$=\frac{1}{n}[\sum\limits_{x\in\mathscr{X}^{0}}q(KQ_{n}/I_{\hat{x}})+\sum\limits_{x\in\mathscr{X}^{1}}q(KQ_{n}/I_{\hat{x}})+\sum\limits_{x\in\mathscr{X}^{2}}q(KQ_{n}/I_{\hat{x}})]$
$=\frac{1}{n}[\sum\limits_{x\in\mathscr{X}^{0}}q(KQ_{n}/I)+\sum\limits_{x\in\mathscr{X}^{1}}q(KQ_{n}/I_{\hat{x}})+\sum\limits_{x\in\mathscr{X}^{2}}q(KQ_{n}/I_{\hat{x}})].$
The following result is a direct consequence of Formula (2).
Corollary 4.4.
Let $A=KQ_{n}/<g>$ be a Nakayama algebra with one generator $g$. Then
$q(A)=\left\{\begin{array}[]{ll}\frac{2}{\ell_{g}}n!,&\ell_{g}\leq n;\\ (n-1)!,&\ell_{g}=n+1;\\ 0,&\ell_{g}>n+1.\end{array}\right.$ (6)
Let $A=KQ_{n}/I$ be a Nakayama algebra with $I=<g_{1},\dots,g_{k}>$. The Formula (2) provides an iteration to compute $q(A)$, as explained in the following example.
Example 4.5.
Let $A=KA_{5}/<g_{1},g_{2}>$ with $g_{1}=\alpha_{2}\alpha_{1}$, $g_{2}=\alpha_{3}\alpha_{2}$ be the algebra in Example 2.3. By Theorem 4.3, we know
$q(A)$ $=\frac{1}{5}[q(A)+\sum\limits_{x\in\mathscr{X}^{1}}q(KQ_{n}/I_{\hat{x}})]$
$=\frac{1}{5}[q(A)+q(KQ_{n}/I_{\hat{1}})+q(KQ_{n}/I_{\hat{4}})]$
$=\frac{1}{4}[q(KQ_{n}/I_{\hat{1}})+q(KQ_{n}/I_{\hat{4}})]$
$=\frac{1}{4}[q(KQ_{n}/<g_{2}>)+q(KQ_{n}/<g_{1}>]$
$=\frac{1}{4}(\frac{2}{3}n!+\frac{2}{3}n!)=\frac{1}{3}n!$
The second to last step uses the result of Corollary 4.4.
4.2. Global supremum and $q$-conjecture
In this subsection, we will give the maximum of $q(A)$ for all Nakayama algebras and prove the $q$-conjecture in this case.
Lemma 4.6.
Let $KQ_{n}/I$ be a quasi-hereditary Nakayama Algebra. If $g$ is a path of $KQ_{n}$ and $g\notin I$, then $q(KQ_{n}/I)>q(KQ_{n}/(I+<g>))$.
Proof.
By Theorem 3.3, $\mathscr{O}(KQ_{n}/I)\supseteq\mathscr{O}(KQ_{n}/(I+<g>))$. Since $g\notin I$, there exists an ordering $\preceq\in\mathscr{O}(KQ_{n}/I)$ such that $h(g)\preceq h(g)+1$ and $d(g)\preceq h(g)+1$. Hence, $\max Hod(g)\in Int(g)$ under $\preceq$. This shows $\preceq\notin\mathscr{O}(KQ_{n}/(I+<g>))$. Therefore, $\mathscr{O}(KQ_{n}/I)\supsetneqq\mathscr{O}(KQ_{n}/(I+<g>))$ and $q(KQ_{n}/I)>q(KQ_{n}/(I+<g>))$. ∎
Theorem 4.7.
Let $A=KQ_{n}/I$ be a Nakayama algebra with I an admissible ideal of $KQ_{n}$. Then
$q(A)\leq\frac{2}{3}n!$ (7)
The equality holds if and only if $I$ is a principal ideal generated by a path of length 3. In particular, the q-conjecture is true for all Nakayama algebras.
Proof.
By Lemma 4.6, $q(KQ_{n}/I)>q(KQ_{n}/I+<g>)$, $g\notin I$. Therefore, $q(A)$ reaches its maximum unless $I$ is a principal ideal. By Corollary 4.4, $\max\{q(KQ_{n}/<g>\}=\frac{2}{3}n!$. ∎
Theorem 4.7 shows that the $q$-conjecture holds for all Nakayama algebras. Unfortunately, the following example shows that the conjecture fails in the case that the involved algebra is not monomial.
Example 4.8.
Let $A=KQ_{n}/I$, $I=<\alpha_{2}\alpha_{1}-\alpha_{4}\alpha_{3}>$
[…]
$q(A)=20>\frac{2}{3}\cdot 4!$.
Declarations
Conflict of interest
The authors state that there is no conflict of interest.
Funding
This research is supported by National Natural Science Foundation of China under Grant No. 123B1026.
Ethical approval
This article does not contain any studies with human participants or animals performed by any of the authors.
Informed consent
Not applicable.
Author Contributions
Both authors contributed equally to this work.
Data Availability Statement
This article does not involve any data, as it is a purely theoretical study.